OpenStudy (hope210):

Will medal and Fan!!!!! Find the solution set of |2x-3| + 1 = 6. A. x = 1 or x = -1 B. x = 5 or x = -2 C. x = 4 or x = -1 D. x = 4 or x = -4

OpenStudy (hope210):

I think it is B but not sure.

OpenStudy (anonymous):

C

OpenStudy (alexandervonhumboldt2):

make it into 2 equations: 2x-3+1=6 -2x+3+1=6 solve

OpenStudy (solomonzelman):

When you have something like \(\color{#000000 }{ \displaystyle \left|x\right|=\ell }\) Then, we are looking for a point(s) on the number line, the distance from which to \(0\) is \(\color{#000000 }{ \displaystyle x }\)

OpenStudy (solomonzelman):